An electron is traveling at 99.9 percent of the velocity of light. What is its mass?

1
Expert's answer
2022-02-13T12:12:37-0500

"m=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}}=2.04\cdot 10^{-29}~kg."